[Hallc_running] [New Logentry] RC Daily Update

dhamilto at jlab.org dhamilto at jlab.org
Sun Dec 2 09:35:01 EST 2018

Logentry Text:

- Data-taking continues to go well, although there wasn't much beam overnight due to a problem with shunts MXU6S03 & MXV4R04, which needed to be replaced.
- Physics analysis yesterday revealed significant suppression of the kaon yield at beam currents of 24 uA and 32 uA relative to 16 uA. The preliminary conclusion is that this is due to blocking of the prompt kaons by accidental pions (see https://logbooks.jlab.org/entry/3633893). We therefore ran overnight at 20 uA. A quick analysis of these runs suggest that this effect might still be present, albeit to a significantly lesser degree, so we are now running at 18 uA.
- Dave M came in to check the issue with BCM4C and the Unser and saw no issue with the latter. He is planning on checking out the BCM4C signal on Monday when the beam is off (https://logbooks.jlab.org/entry/3633906).

RC Meeting:
- There will be no RC meeting today, although I will be around the counting house all afternoon if anyone wants to discuss anything.

Notices in effect for shift crews:
- Shift leaders, please fill out the check list once per shift.
- Shift leaders, please do a quick cross-check of the BTA numbers against your shift summary.
- Check that the pre-scaler gui is active and hasn't frozen / doesn't need re-started.
- Be on the lookout for any errors when running the replay, particularly errors concerning missing reference times.
- Note the new beam positions for this energy (on the whiteboard / note on wall)
- NB If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''

Planned accelerator configuration changes:
- Beam studies / system recovery planned for Monday, Day shift.

Planned accesses to the hall on Monday (during beam studies):
- Mark and Ethan plan to reboot the HMS collimator motion controller.
- Joe will look at possibly changing the A/C filter on the HMS shield hut..
- If there is anything else, please let me know.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3634046
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181202/fa9616de/attachment.html>

More information about the Hallc_running mailing list